Break-Even Sales Anheuser-Busch InBev Companies, Inc., reported the following operating information for a recent year (in millions of dollars): Net sales $39,758 Cost of goods sold $16,447 Selling, general and administration 10,578 $27,025 Income from operations $12,733* *Before special items In addition, assume that Anheuser-Busch InBev sold 320 million barrels of beer during the year. Assume that variable costs were 70% of the cost of goods sold and 40% of selling, general, and administration expenses. Assume that the remaining costs are fixed. For the following year, assume that Anheuser-Busch InBev expects pricing, variable costs per barrel, and fixed costs to remain constant, except that new distribution and general office facilities are expected to increase fixed costs by $400 million. When computing the cost per unit amounts for the break-even formula, round to two decimal places. If required, round your final answer to the nearest whole barrel.

a. Compute the break-even number of barrels for the current year. ________barrels

b. Compute the anticipated break-even number of barrels for the following year. ________barrels
Answer #1

SOLUTION

1. Sales price per barrel = $39,758 / 320 = 124.24

Variable cost per barrel = {Cost of goods sold + Selling, general and administration expenses} / 320

= {(16,447*70%) + (10,578*40%)} / 320

= {11,512.90 + 4,231.20} / 320

= 15,744.10 / 320 = $49.20

Contribution margin = 124.24 - 49.20 = 75.04

Fixed cost = (16,447*30%) + (10,578*60%)

= 4,934.10 + 6,346.80

= 11,280.90

Break even point = Fixed costs / Contribution margin per barrel

= 11,280.90 / 75.04 = 150 barrels

2. Anticipated break-even number = Revised Fixed costs / Contribution margin per barrel

= (11,280.90 + 400) / 75.04

= 11,680.90 / 75.04 = 156